# Blind Inverse Game Theory: Jointly Decoding Rewards and Rationality in Entropy-Regularized Competitive Games

### v1

@@ 1 INTRODUCTION
The statistical inference of underlying objectives from observed strategic behavior is a fundamental challenge spanning artificial intelligence, economics, and multi-agent systems (Ng&Russell2000; Arora&Doshi2021). In multi-agent environments, this problem is formalized as Inverse Game Theory (IGT) (Kuleshov&Schrijvers2015; Yu2019), where the goal is to identify the payoff structures that rationalize observed equilibrium play.
Recovering rewards in competitive settings, particularly two-player zero-sum games, presents significant statistical difficulties. Standard IGT approaches based on the Nash Equilibrium (NE) (Nash1951) concept are generally ill-posed, leading to fundamental identification issues (Ahuja&Orlin2001; Metelli2021), equilibrium multiplicity, and computational intractability (Wang&Klabjan2018; Wu2022).
To address this primary identification challenge, the literature has increasingly adopted models of bounded rationality, notably through entropy regularization (Mertikopoulos&Sandholm2016). This smooths the agents’ best responses, resulting in a unique equilibrium known as the Quantal Response Equilibrium (QRE) (McKelvey&Palfrey1995). The QRE framework has facilitated improved identifiability in single-agent Inverse Reinforcement Learning (IRL) (Cao2021; Rolland2022) and enabled efficient reward recovery in zero-sum IGT by transforming the problem into a tractable system of linear equations (Liao2025).
@@ 1.1 The Challenge: Unknown Rationality and Scale Ambiguity
The assumption of a known $\tau$ is problematic for statistical inference. In practice, the agents’ rationality is unknown and must be estimated from the data. Critically, if $\tau$ is unknown, a fundamental identification problem re-emerges. The reward parameters ($\theta$) and the temperature ($\tau$) are inextricably coupled in the QRE definition, which depends only on the ratio $Q/\tau$. This results in an inherent multiplicative scale ambiguity: if a pair $(\theta,\tau)$ explains the observed behavior, then any scaled pair $(k\theta,k\tau)$ for $k>0$ provides an identical explanation. This ambiguity constitutes a fundamental roadblock because the true parameters become statistically unidentifiable. Intuitively, we cannot distinguish between a highly rational agent (small $\tau$) playing a low-stakes game (small $\theta$) and a highly irrational agent (large $\tau$) playing a high-stakes game (large $\theta$).
Consequently, an incorrect specification of $\tau$ leads to a systematic bias in the scale of the recovered rewards, leading to model misspecification and invalidating theoretical guarantees.
@@ 1.2 Our Contributions
To address this fundamental gap, we introduce Blind Inverse Game Theory (Blind-IGT), the first statistical framework enabling the simultaneous recovery of reward parameters ($\theta$) and the unknown temperature ($\tau$) from observed equilibrium behavior in entropy-regularized zero-sum games. This requires solving a challenging bilinear inverse problem. By introducing a normalization constraint on the rewards, we resolve the inherent scale ambiguity and provide a comprehensive theoretical analysis. Our main contributions are:
  • Comprehensive Identifiability Analysis: We formally characterize the multiplicative scale ambiguity. We establish necessary and sufficient conditions for unique joint identification (Theorem 1). Furthermore, we provide a rigorous framework for partial identification via confidence sets when these conditions fail (Proposition 1).
  • Efficient NLS Estimator with Optimal Rates: We propose a novel Normalized Least Squares (NLS) estimator (Algorithm 1). We provide rigorous finite-sample guarantees, proving NLS achieves the optimal parametric convergence rate of $\mathcal{O}(N^{-1/2})$ for the joint recovery problem (Theorem 2).
  • Robust Extension to Markov Games: We extend our framework to dynamic settings (Theorem 3). Furthermore, we empirically demonstrate that NLS tracks the optimal convergence rates even when transition dynamics are unknown and estimated from data (Section 6.4).
  • Empirical Validation: Extensive simulations confirm the statistical efficiency of NLS, its necessity when $\tau$ is unknown, and its robustness to misspecification.
@@ 2 RELATED WORK
Inverse Optimization (IO) and Inverse Reinforcement Learning (IRL). IO aims to infer objective functions from observed decisions (Ahuja&Orlin2001; Chan2022). IRL specifically focuses on recovering reward functions in sequential decision-making (Ng&Russell2000). A prominent approach is Maximum Entropy IRL (MaxEnt-IRL) (Ziebart2008; Wulfmeier2016; Finn2016).
Recent theoretical work has analyzed reward identifiability in entropy-regularized settings. Cao2021 and Rolland2022 demonstrated improved identifiability in single-agent scenarios. Crucially, this prior work addresses an additive ambiguity inherent in the Bellman equation (potential shaping). In contrast, our work tackles a multiplicative scale ambiguity arising from the unknown temperature in a competitive setting. These represent distinct mathematical challenges.
Inverse Game Theory (IGT). IGT extends IRL to multi-agent systems, aiming to recover the underlying payoff structures that rationalize observed strategic interactions (Vorobeychik2007; Arora&Doshi2021). The predominant approach in the literature assumes that the observed behavior corresponds to a standard equilibrium concept, typically the Nash Equilibrium (NE) (Lin2014; Kuleshov&Schrijvers2015) or the Correlated Equilibrium (CE) (Waugh2011). However, these approaches often struggle with non-uniqueness and computational intractability, frequently relying on complex bilevel optimization as discussed in existing work (Wang&Klabjan2018; Wu2022; Konstantakopoulos2017).
Entropy Regularization and QRE. The Quantal Response Equilibrium (QRE) (McKelvey&Palfrey1995) models agents playing noisy best responses. Unlike the idealized Nash Equilibrium (NE), QRE provides smooth, unique equilibrium predictions, which facilitates tractable analysis (Mertikopoulos&Sandholm2016) and often better fits empirical data from human behavior (Goeree2016; Camerer2003). The temperature parameter $\tau$ explicitly controls the degree of bounded rationality, allowing QRE to interpolate between uniform random play (as $\tau\to\infty$) and the NE (as $\tau\to 0$). This connection makes QRE a central concept in modern multi-agent reinforcement learning (MARL), where entropy regularization is widely employed to stabilize training and encourage exploration (Haarnoja2018; Cen2021; Guan2021; Ahmed2019; Zhan2023).
@@
The work most closely related to ours is Liao2025, which established a framework for IGT in entropy-regularized zero-sum games based on the linearized QRE constraints we utilize. Our work directly addresses the fundamental limitation of their analysis (the assumption of a known $\tau$) by developing a ”blind” framework. We provide a methodology (NLS) for unique point identification of both rewards and the unknown temperature via normalization. Furthermore, we extend the partial identification analysis of Liao2025 to the bilinear setting where $\tau$ is unknown.
Bilinear Inverse Problems. The structure of Blind-IGT is a bilinear inverse problem, where we seek two vectors whose product explains the observations. Such problems arise in areas like blind deconvolution (Ahmed2014) and self-calibration (Ling&Strohmer2015). They are generally non-convex. Our approach leverages the specific structure of the QRE constraints and the normalization constraint to derive an efficient and provably correct estimation method.
@@ 3.1 Matrix Games
Entropy Regularization. We study the game under entropy regularization (Mertikopoulos&Sandholm2016). The regularized game is formulated as:
$\max_{\mu\in\Delta(\mathcal{A})}\min_{\nu\in\Delta(\mathcal{B})}\left\{\mu^{\top}Q\nu+\tau\mathcal{H}(\mu)-\tau\mathcal{H}(\nu)\right\},$ (1)
where $\mu\in\Delta(\mathcal{A})$ and $\nu\in\Delta(\mathcal{B})$ are the mixed strategies for each player, $\tau>0$ is the regularization parameter (temperature), and $\mathcal{H}(\pi)=-\sum_{i}\pi_{i}\log(\pi_{i})$ is the Shannon entropy (Shannon1948a).
Quantal Response Equilibrium (QRE). The optimization problem (1) admits a unique solution $(\mu^{*},\nu^{*})$, known as the QRE (McKelvey&Palfrey1995). The QRE is characterized by the following fixed-point equations (logit responses):
$\mu^{*}(a)$ $=\frac{\exp(Q(a,\cdot)\nu^{*}/\tau)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(Q(a^{\prime},\cdot)\nu^{*}/\tau)},$ (2)
$\nu^{*}(b)$ $=\frac{\exp(-Q(\cdot,b)^{\top}\mu^{*}/\tau)}{\sum_{b^{\prime}\in\mathcal{B}}\exp(-Q(\cdot,b^{\prime})^{\top}\mu^{*}/\tau)}.$ (3)
These non-linear equations can be linearized by taking the logarithm and normalizing with respect to a reference action (e.g., action 1). This transformation (detailed in Appendix B.1) yields the following system of m+n2m+n-2 linear constraints:
$\begin{cases}(Q(a,\cdot)-Q(1,\cdot))\nu^{*}=\tau\cdot\log\frac{\mu^{*}(a)}{\mu^{*}(1)},&\forall a\in\mathcal{A}\setminus\{1\}\\ (Q(\cdot,1)-Q(\cdot,b))^{\top}\mu^{*}=\tau\cdot\log\frac{\nu^{*}(b)}{\nu^{*}(1)},&\forall b\in\mathcal{B}\setminus\{1\}\end{cases}.$ (4)
@@ 3.2 Markov Games
$Q^{\mu,\nu}(s,a,b)$ $=r(s,a,b)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a,b)}[V^{\mu,\nu}(s^{\prime})],$ (5)
@@
$\quad+\tau\mathcal{H}(\mu(s))-\tau\mathcal{H}(\nu(s)).$ (6)
The QRE $(\mu^{*},\nu^{*})$ is the unique solution to the minimax problem $V^{*}(s)=\max_{\mu}\min_{\nu}V^{\mu,\nu}(s)$. Similar to matrix games, the QRE policies at each state $s$ satisfy the fixed-point equations defined by (2) and (3), using the Q-function $Q^{*}(s)$ in place of the payoff matrix $Q$.
@@ 4.1 Problem Formulation and Ambiguity
Assumption 1 (Linear Payoff Functions).
@@
Under Assumption 1, the linearized QRE constraints in (4) can be rewritten in a compact matrix-vector form. We define the feature matrices $A(\nu)\in\mathbb{R}^{(m-1)\times d}$ and $B(\mu)\in\mathbb{R}^{(n-1)\times d}$, and the log-ratio vectors $c(\mu)\in\mathbb{R}^{m-1}$ and $d(\nu)\in\mathbb{R}^{n-1}$.
@@
$X(\mu^{*},\nu^{*})\theta^{*}=\tau^{*}\cdot y(\mu^{*},\nu^{*}).$ (7)
The Scale Ambiguity. The system (7) highlights the fundamental identification challenge. If $(\theta^{*},\tau^{*})$ is a solution, then any scaled pair $(k\theta^{*},k\tau^{*})$ for $k>0$ is also a solution that induces the exact same QRE $(\mu^{*},\nu^{*})$.
@@ 4.2 Identifiability Conditions
Assumption 2 (Normalization Constraint).
@@
Assumption 2 is necessary for the exact recovery of the scales of both $\theta^{*}$ and $\tau^{*}$.
Remark 1 (Robustness to Misspecification).
@@
Theorem 1 (Identifiability of Blind-IGT).
Under Assumptions 1 and 2 (with $C=1$), the pair $(\theta^{*},\tau^{*})$ is uniquely identifiable from the QRE $(\mu^{*},\nu^{*})$ if and only if the following two conditions hold:
@@
Proof Sketch (Appendix C.1).
Sufficiency relies on the Rank Condition ensuring a unique direction for $\theta^{*}$ via the Moore-Penrose inverse (Penrose1955). The Normalization Constraint then uniquely determines the scale (and thus $\tau^{*}$), provided the Non-Uniformity Condition holds (ensuring $\theta^{*}\neq 0$). Necessity is shown by constructing counterexamples when either condition fails. ∎
@@ 4.3 Estimation via Normalized Least Squares
We propose the Normalized Least Squares (NLS) estimator (Algorithm 1). The core idea is to first estimate the ratio $\theta^{*}/\tau^{*}$ using standard least squares, and then use the normalization constraint to disentangle the two parameters.
Algorithm 1 Normalized Least Squares (NLS) for Blind-IGT
1:Input: Data $\{(a^{k},b^{k})\}_{k=1}^{N}$, features $\phi$, Norm $C$.
[…]
4:Compute the standard least squares solution:
[…]
6:Estimate the temperature:
[…]
8:Estimate the reward parameters:
[…]

### v2

@@ 1 Introduction
The statistical inference of underlying objectives from observed strategic behavior is a fundamental challenge spanning artificial intelligence, economics, and multi-agent systems [Ng and Russell, 2000, Arora and Doshi, 2021]. In multi-agent environments, this problem is formalized as Inverse Game Theory (IGT) [Kuleshov and Schrijvers, 2015, Yu et al., 2019], where the goal is to identify the payoff structures that rationalize observed equilibrium play.
Recovering rewards in competitive settings, particularly two-player zero-sum games, presents significant statistical difficulties. Standard IGT approaches based on the Nash Equilibrium (NE) [Nash, 1951] concept are generally ill-posed, leading to fundamental identification issues [Ahuja and Orlin, 2001, Metelli et al., 2021], equilibrium multiplicity, and computational intractability [Wang and Klabjan, 2018, Wu et al., 2022].
To address this primary identification challenge, the literature has increasingly adopted models of bounded rationality, notably through entropy regularization [Mertikopoulos and Sandholm, 2016]. This smooths the agents’ best responses, resulting in a unique equilibrium known as the Quantal Response Equilibrium (QRE) [McKelvey and Palfrey, 1995]. The QRE framework has facilitated improved identifiability in single-agent Inverse Reinforcement Learning (IRL) [Cao et al., 2021, Rolland et al., 2022] and enabled efficient reward recovery in zero-sum IGT by transforming the problem into a tractable system of linear equations [Liao et al., 2025].
@@ 1.1 The Challenge: Unknown Rationality and Scale Ambiguity
The assumption of a known $\tau$ is problematic for statistical inference. In practice, the agents’ rationality is unknown and must be estimated from the data. Critically, if $\tau$ is unknown, a fundamental identification problem re-emerges. The reward parameters ($\theta$) and the temperature ($\tau$) are inextricably coupled in the QRE definition, which depends only on the ratio $Q/\tau$. This results in an inherent multiplicative scale ambiguity: if a pair $(\theta,\tau)$ explains the observed behavior, then any scaled pair $(k\theta,k\tau)$ for $k>0$ provides an identical explanation. This ambiguity constitutes a fundamental roadblock because the true parameters become statistically unidentifiable. Intuitively, we cannot distinguish between a highly rational agent (small $\tau$) playing a low-stakes game (small $\theta$) and a highly irrational agent (large $\tau$) playing a high-stakes game (large $\theta$). Consequently, an incorrect specification of $\tau$ leads to a systematic bias in the scale of the recovered rewards, leading to model misspecification and invalidating theoretical guarantees.
@@ 1.2 Our Contributions
To address this fundamental gap, we introduce Blind Inverse Game Theory (Blind-IGT), the first statistical framework enabling the simultaneous recovery of reward parameters ($\theta$) and the unknown temperature ($\tau$) from observed equilibrium behavior in entropy-regularized zero-sum games. This requires solving a challenging bilinear inverse problem. By introducing a normalization constraint on the rewards, we resolve the inherent scale ambiguity and provide a comprehensive theoretical analysis. Our main contributions are as follows.
First, we provide a comprehensive identifiability analysis where we formally characterize the multiplicative scale ambiguity, establish necessary and sufficient conditions for unique joint identification (Theorem 4.4), and provide a rigorous framework for partial identification via confidence sets when these conditions fail (Proposition 4.8).
Second, we propose an efficient Normalized Least Squares (NLS) estimator with optimal rates, providing rigorous finite-sample guarantees that prove NLS achieves the optimal parametric convergence rate of $\mathcal{O}(N^{-1/2})$ for the joint recovery problem (Theorem 4.6).
Third, we develop a robust extension to Markov games (Theorem 5.4) and empirically demonstrate that NLS tracks the optimal convergence rates even when transition dynamics are unknown and estimated from data (Section 6.4).
Finally, extensive simulations confirm the statistical efficiency of NLS, its necessity when $\tau$ is unknown, and its robustness to misspecification.
@@ 2 Related Work
Inverse Optimization (IO) and Inverse Reinforcement Learning (IRL). IO aims to infer objective functions from observed decisions [Ahuja and Orlin, 2001, Chan et al., 2023]. IRL specifically focuses on recovering reward functions in sequential decision-making [Ng and Russell, 2000]. A prominent approach is Maximum Entropy IRL (MaxEnt-IRL) [Ziebart et al., 2008, Wulfmeier et al., 2016, Finn et al., 2016].
Recent theoretical work has analyzed reward identifiability in entropy-regularized settings. Cao et al. [2021] and Rolland et al. [2022] demonstrated improved identifiability in single-agent scenarios. Crucially, this prior work addresses an additive ambiguity inherent in the Bellman equation (potential shaping). In contrast, our work tackles a multiplicative scale ambiguity arising from the unknown temperature in a competitive setting. These represent distinct mathematical challenges.
Inverse Game Theory (IGT). IGT extends IRL to multi-agent systems, aiming to recover the underlying payoff structures that rationalize observed strategic interactions [Vorobeychik et al., 2007, Arora and Doshi, 2021]. The predominant approach in the literature assumes that the observed behavior corresponds to a standard equilibrium concept, typically the Nash Equilibrium (NE) [Lin et al., 2018, Kuleshov and Schrijvers, 2015] or the Correlated Equilibrium (CE) [Waugh et al., 2011]. However, these approaches often struggle with non-uniqueness and computational intractability, frequently relying on complex bilevel optimization as discussed in existing work [Wang and Klabjan, 2018, Wu et al., 2022, Konstantakopoulos et al., 2018].
Entropy Regularization and QRE. The Quantal Response Equilibrium (QRE) [McKelvey and Palfrey, 1995] models agents playing noisy best responses. Unlike the idealized Nash Equilibrium (NE), QRE provides smooth, unique equilibrium predictions, which facilitates tractable analysis [Mertikopoulos and Sandholm, 2016] and often better fits empirical data from human behavior [Goeree et al., 2016, Camerer, 2003]. The temperature parameter $\tau$ explicitly controls the degree of bounded rationality, allowing QRE to interpolate between uniform random play (as $\tau\to\infty$) and the NE (as $\tau\to 0$). This connection makes QRE a central concept in modern multi-agent reinforcement learning (MARL), where entropy regularization is widely employed to stabilize training and encourage exploration [Haarnoja et al., 2018, Cen et al., 2021, Guan et al., 2021, Ahmed et al., 2019, Zhan et al., 2023].
@@
The work most closely related to ours is Liao et al. [2025], which established a framework for IGT in entropy-regularized zero-sum games based on the linearized QRE constraints we utilize. Our work directly addresses the fundamental limitation of their analysis (the assumption of a known $\tau$) by developing a "blind" framework. We provide a methodology (NLS) for unique point identification of both rewards and the unknown temperature via normalization. Furthermore, we extend the partial identification analysis of Liao et al. [2025] to the bilinear setting where $\tau$ is unknown.
Bilinear Inverse Problems. The structure of Blind-IGT is a bilinear inverse problem, where we seek two vectors whose product explains the observations. Such problems arise in areas like blind deconvolution [Ahmed et al., 2014] and self-calibration [Ling and Strohmer, 2015]. They are generally non-convex. Our approach leverages the specific structure of the QRE constraints and the normalization constraint to derive an efficient and provably correct estimation method.
@@ 3.1 Matrix Games
Entropy Regularization. We study the game under entropy regularization [Mertikopoulos and Sandholm, 2016]. The regularized game is formulated as:
$\max_{\mu\in\Delta(\mathcal{A})}\min_{\nu\in\Delta(\mathcal{B})}\left\{\mu^{\top}Q\nu+\tau\mathcal{H}(\mu)-\tau\mathcal{H}(\nu)\right\},$ (3.1)
where $\mu\in\Delta(\mathcal{A})$ and $\nu\in\Delta(\mathcal{B})$ are the mixed strategies for each player, $\tau>0$ is the regularization parameter (temperature), and $\mathcal{H}(\pi)=-\sum_{i}\pi_{i}\log(\pi_{i})$ is the Shannon entropy [Shannon, 1948].
Quantal Response Equilibrium (QRE). The optimization problem (3.1) admits a unique solution $(\mu^{*},\nu^{*})$, known as the QRE [McKelvey and Palfrey, 1995]. The QRE is characterized by the following fixed-point equations (logit responses):
$\mu^{*}(a)$ $=\frac{\exp(Q(a,\cdot)\nu^{*}/\tau)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(Q(a^{\prime},\cdot)\nu^{*}/\tau)},$ (3.2)
$\nu^{*}(b)$ $=\frac{\exp(-Q(\cdot,b)^{\top}\mu^{*}/\tau)}{\sum_{b^{\prime}\in\mathcal{B}}\exp(-Q(\cdot,b^{\prime})^{\top}\mu^{*}/\tau)}.$ (3.3)
These non-linear equations can be linearized by taking the logarithm and normalizing with respect to a reference action (e.g., action 1). This transformation (detailed in Appendix C.1) yields the following system of m+n2m+n-2 linear constraints:
$\begin{cases}(Q(a,\cdot)-Q(1,\cdot))\nu^{*}=\tau\cdot\log\frac{\mu^{*}(a)}{\mu^{*}(1)},&\forall a\in\mathcal{A}\setminus\{1\}\\ (Q(\cdot,1)-Q(\cdot,b))^{\top}\mu^{*}=\tau\cdot\log\frac{\nu^{*}(b)}{\nu^{*}(1)},&\forall b\in\mathcal{B}\setminus\{1\}\end{cases}.$ (3.4)
@@ 3.2 Markov Games
$Q^{\mu,\nu}(s,a,b)$ $=r(s,a,b)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a,b)}[V^{\mu,\nu}(s^{\prime})],$ (3.5)
@@
$\quad+\tau\mathcal{H}(\mu(s))-\tau\mathcal{H}(\nu(s)).$ (3.6)
The QRE $(\mu^{*},\nu^{*})$ is the unique solution to the minimax problem $V^{*}(s)=\max_{\mu}\min_{\nu}V^{\mu,\nu}(s)$. Similar to matrix games, the QRE policies at each state $s$ satisfy the fixed-point equations defined by (3.2) and (3.3), using the Q-function $Q^{*}(s)$ in place of the payoff matrix $Q$.
@@ 4.1 Problem Formulation and Ambiguity
Assumption 4.1 (Linear Payoff Functions).
@@
Under Assumption 4.1, the linearized QRE constraints in (3.4) can be rewritten in a compact matrix-vector form. We define the feature matrices $A(\nu)\in\mathbb{R}^{(m-1)\times d}$ and $B(\mu)\in\mathbb{R}^{(n-1)\times d}$, and the log-ratio vectors $c(\mu)\in\mathbb{R}^{m-1}$ and $d(\nu)\in\mathbb{R}^{n-1}$.
@@
$X(\mu^{*},\nu^{*})\theta^{*}=\tau^{*}\cdot y(\mu^{*},\nu^{*}).$ (4.1)
The Scale Ambiguity. The system (4.1) highlights the fundamental identification challenge. If $(\theta^{*},\tau^{*})$ is a solution, then any scaled pair $(k\theta^{*},k\tau^{*})$ for $k>0$ is also a solution that induces the exact same QRE $(\mu^{*},\nu^{*})$.
@@ 4.2 Identifiability Conditions
Assumption 4.2 (Normalization Constraint).
@@
Assumption 4.2 is necessary for the exact recovery of the scales of both $\theta^{*}$ and $\tau^{*}$.
Remark 4.3 (Robustness to Misspecification).
@@
Theorem 4.4 (Identifiability of Blind-IGT).
Under Assumptions 4.1 and 4.2 (with $C=1$), the pair $(\theta^{*},\tau^{*})$ is uniquely identifiable from the QRE $(\mu^{*},\nu^{*})$ if and only if the following two conditions hold:
@@
Proof Sketch (Appendix D.1).
Sufficiency relies on the Rank Condition ensuring a unique direction for $\theta^{*}$ via the Moore-Penrose inverse [Penrose, 1955]. The Normalization Constraint then uniquely determines the scale (and thus $\tau^{*}$), provided the Non-Uniformity Condition holds (ensuring $\theta^{*}\neq 0$). Necessity is shown by constructing counterexamples when either condition fails. ∎
@@ 4.3 Estimation via Normalized Least Squares
We propose the Normalized Least Squares (NLS) estimator. The complete pseudocode is provided in Appendix A. The core idea is to first estimate the ratio $\theta^{*}/\tau^{*}$ using standard least squares, and then use the normalization constraint to disentangle the two parameters.